# Spectroscopy using a visible photonic lantern at the Subaru telescope: Laboratory characterization and first on-sky demonstration on Ikiiki ( Leo) and ‘Aua ( Ori)

### v1

@@ 4.1 Observation log
The focal ratio selected for this observation was 8, providing the best injection efficiency. We observed Ikiiki ( Leo, = 1.37) and ‘Aua ( Ori, = -1.17) on February 18th 2024 UTC during an Engineering night (Proposal ID: S24A-EN01, PI : Julien Lozi, Support Astronomers: Julien Lozi and Sébastien Vievard, Telescope Operator : Erin Dailey). Data was acquired at a 200 Hz framerate (see Table 1). The seeing in H-band was estimated to be about ″ by the SCExAO NIR internal fast frame camera. The VAMPIRES focal plane camera was used to estimate the residual tip/tilt and the Strehl in the visible, during the observation of Ikiiki, at 670 nm, 720 nm and 760 nm. The tip/tilt residual was about 12 mas RMS across all wavelengths. The estimated Strehl ration was on average of at 670 nm, at 720 nm, and at 760 nm.
@@ 4.2 Observation of Ikiiki
We observed Ikiiki for 20 minutes at a 200 Hz frame rate, providing a total of 240,000 frames. Because the observing conditions were not particularly good, all the acquired frames could not be exploited. We selected the best of the frames based on the flux received by the PL science camera. Figure 14 shows a dark-subtracted averaged image after frame selection. We can already identify, in this data, three prominent absorption lines. We also observe that the flux distribution varies across traces, indicating the sensitivity of PL to the input scene/wavefront. We calculated the Signal-to-Noise Ratio (SNR) for each minute of observation by dividing the average flux by the standard deviation of the flux, adjusted for the number of frames (i.e., the error on the flux measurement), per spectral channel over time. This approach yielded SNR values that varied around 200, primarily due to fluctuations in injection efficiency.
@@ 4.2.1 Spectrum reconstruction
For each selected frame, we subtract the dark+bias and extract each PL output trace. The traces are then co-added for each frame, and then averaged over the entire observation sequence.
To calibrate the response of our instrument, we use a flat field halogen lamp (approximately black body spectrum) positioned at the entrance of AO188 to uniformly illuminate the PL input FoV. By performing the same computation as previously described, we derive a flat field spectrum, which we then calibrate using the theoretical black body slope corresponding to the Halogen lamp temperature (3020.3K). Subsequently, we utilize the resulting calibrated flat data to calibrate the Ikiiki reconstructed spectrum obtained with the PL. We further refine the computed spectrum by normalizing it to match a reference spectrum obtained from Alekseeva et al. (1996), represented by black dots in Figure 15.
On the same graph, we show the resulting calibrated and adjusted spectrum of Ikiiki obtained with the PL. The latter closely matches the reference spectrum. Some differences can be noted, which may be attributed to the imperfect calibration of our data. Indeed, calibrating our reconstructed spectrum with the halogen lamp may potentially overlook chromatic perturbations caused by the atmosphere and/or various optical surfaces prior to AO188. Ideally, calibration should be conducted using a well-calibrated point source in the sky. The spectrum of Ikiiki exhibits three distinct absorption lines: Atmospheric oxygen A and B bands (around 761 nm and 687 nm, respectively), and H (at approximately 656 nm). The absorption in the two Oxygen bands is due to Oxygen molecules in Earth’s atmosphere and does not appear in the reference spectrum due to its terrestrial origin. At the R3000 spectral resolution, the absorption bands appear as a deep broad absorption line with and a forest of narrower lines at longer wavelength. The H absorption line originates from the cooler outer layers of the Ikiiki atmosphere, where neutral hydrogen atoms absorb photons emitted by the deeper, hotter layers at the energy level corresponding to the H transition.
@@ 4.2.2 Injection efficiency
In order to estimate the on-sky injection efficiency of the PL, we compared the acquired spectra (prior to flat-field calibration) to the reference spectro-photometric data of Ikiiki (Alekseeva et al., 1996). We summarize in Table 2 the throughput of elements from the entrance of the telescope down to the detector. We will consider the detector quantum efficiency provided by the manufacturer (from to from 600 nm to 785 nm). We also modeled the throughput of the Earth’s atmosphere using the synthetic photometry python package (synphot) maintained by the Space Telescope Science Institute (STScI). The airmass during the observation was 1.01, and induced an extinction ranging from 0.1 to 0.03 magnitude between 600 nm and 770 nm, leading to a throughput of about in average.
@@
Even under these difficult conditions, the collected data had a good SNR and allowed for a faithful reconstruction of the Ikiiki spectrum.
@@ 4.3 Observation of ‘Aua
We observed ‘Aua for 53 minutes at a 200 Hz frame rate, providing a total of 636,000 frames. Similarly to the previous target, we performed a frame selection since the conditions were not good. The best frames were selected, dark+bias subtracted and averaged to compute the image displayed on Figure 18. Many spectral features can be identified on the averaged image. The computed SNR per minute was around 130, a sign that the observation conditions worsened compared to the previous target (where the SNR per minute was 200 in average).
After trace extraction and calibration with flat-field data, we co-added the resulting spectra to obtain the spectrum shown in Figure 19. The spectrum shows multiple absorption spectral lines. The absorption lines detected on Ikiiki are present: Oxygen bands A and B from the Earth’s atmosphere around 761 nm and 687 nm respectively, and H absorption by the stellar atmosphere at about 656 nm. Many titanium oxide (TiO) lines are also present in the spectrum, with the strongest at about 705 nm. TiO lines in red supergiants are characteristic of convection cells within the star’s atmosphere (Young et al., 2003; Afram & Berdyugina, 2019), and are well-known in the case of ’Aua as they are extensively exploited for image reconstruction using interferometry (Young et al., 2000; Levesque et al., 2005; Chiavassa et al., 2010). The ability to detect these (and the following) spectral lines while conserving spatial information via use of a PL opens the door to future use of differential phase interferometry (Akeson et al., 2000) to recover spatial features (such as stellar surface details) from PL observations. We can also spot the presence of vanadium oxide lines (VO) from about 730 nm to 750 nm. Similarly to TiO lines, VO lines are generated by convection cells in red supergiant stars, although they are less dominant than TiO. Finally we identified the Barium II (Ba II) line at about 649 nm (Koza, 2010). These lines can also be seen in data from the X-shooter Spectral Library (XSL) (Gonneau et al., 2020) taken at the VLT-UT2 using a medium resolution spectrograph. This demonstration highlights the capabilities of the PL signal reconstruction to retrieve a target’s spectral features. The next step, not addressed in this demonstration paper, will be to exploit the spectral reconstruction to extract astronomical information about the observed target.
@@ 5 Conclusion
We presented the integration, characterization and on-sky demonstration of a photonic lantern at the 8.2 m Subaru telescope. Photonic lanterns allow for high efficiency SMF-based spectroscopy thanks to their unique design. Their MM input insures high coupling efficiency, compared to a single SMF, especially at visible wavelengths where even ExAO systems struggle to reach high Strehl ratios. The SM outputs of PLs are perfectly suited for spectroscopy using compact and stable instruments.
@@
Depending on the optical injection setup focal ratio, the PL can provide various field of view and injection efficiency performance depending on the desired science output. Our setup shows an optimal injection efficiency for a focal ratio of about 8, providing an injection efficiency of about over a field of view of 80 mas, with an efficiency peak of at 642 nm.
@@
We studied the relationship between the PL injection efficiency and Strehl ratio. Various turbulence profiles were simulated on the SCExAO DM, and the results were confronted to simulations in the case of an SMF. We showed that PLs have a similar linear relationship between the injection efficiency and the Strehl to that of an SMF. However, PLs show better performance in the presence of turbulence dominated by low frequency aberrations.
@@
Finally, we tested the PL on-sky. Two stars were observed: a point source, Ikiiki ( Leo) and a red supergiant star, ‘Aua ( Ori). We were first able to reconstruct Ikiiki’s spectrum with high fidelity compared to its reference. Under median seeing conditions, the injection efficiency into the PL averaged to and reached up to at 642 nm, leading to an instrument throughput up to . The best performance was obtained at 680 nm, where the injection efficiency into the PL averaged to and reached up to , leading to an instrument throughput up to . It is important to note that these figures likely fall short of demonstrating the full capabilities of our visible PL under more favorable conditions. For context, attempting a similar light injection directly into a single SMF under these conditions would likely yield inferior results. The observation of ‘Aua confirmed our ability to reconstruct a complex spectrum from the PL data. We were able to retrieve multiple spectral features characteristic of this target.
@@
The on-sky performance should be increased under better atmospheric conditions, and even further with the soon-to-be upgrades of the Subaru AO facility DM (from 188 elements to 3000 elements - ) and visible wavefront sensor (Ahn et al., 2023). We also plan on upgrading the NPBS, picking up the light from SCExaO and routing it to the PL, with a R90/T10 beamsplitter. This should significantly increase the overall throughput of the instrument.
@@
This result shows that our setup can be used for high-throughput SMF-fed spectroscopy in the visible, especially for H spectro-imaging. The ongoing and near-futur work aims to provide this capability to the community as a feature of the FIRST instrument. This demonstration constitutes the first on-sky acquisition of a spectrum using a photonic lantern fed by an Extreme Adaptive Optics system in the visible.
@@ Acknowledgements.
- Kim et al. (2024) Kim, Y. J., Fitzgerald, M. P., Lin, J., et al. 2024, arXiv e-prints, arXiv:2402.08158
@@
- Tobin et al. (2024) Tobin, T. L., Currie, T., Li, Y., et al. 2024, arXiv e-prints, arXiv:2403.04000
@@
Appendix A Optical setup parameters for various numerical apertures of the beam injected into the Photonic Lantern
Table 3 provides optimized parameters of the Photonic Lantern injection bench for reference. Parameters are given for various focal ratios of the beam injected into the Photonic Lantern. Distances between the optical components are measured from the back surface of the first optical component to the front surface of the second optical component. For focal ratios of , , and , only two lenses are needed (L1 and L2). The back focal length provided in the table corresponds to the distance between the back surface of the last lens (L2 or L3, depending on the focal ratio considered) and the focal plane.

### v3

@@ 4.1 Observation log
The focal ratio selected for this observation was 8, providing the best injection efficiency. We observed Ikiiki ( Leo, = 1.37) and ‘Aua ( Ori, = -1.17) on February 18, 2024 UTC during an engineering night (Proposal ID: S24A-EN01, PI : Julien Lozi, Support Astronomers: Julien Lozi and Sébastien Vievard, Telescope Operator : Erin Dailey). Data was acquired at a 200 Hz framerate (see Table 1). The seeing in the H band was estimated to be about ″ by the SCExAO NIR internal fast frame camera. The VAMPIRES focal plane camera was used to estimate the residual tip or tilt and the Strehl in the visible, during the observation of Ikiiki, at 670 nm, 720 nm and 760 nm. The tip-tilt residual was about 12 mas RMS across all wavelengths. The estimated Strehl ratio was on average at 670 nm, at 720 nm, and at 760 nm.
@@ 4.2 Observation of Ikiiki
We observed Ikiiki for 20 minutes at a 200 Hz frame rate, providing a total of 240,000 frames. Because the observing conditions were not particularly good, not all of the acquired frames could be exploited. We selected the best of the frames based on the flux received by the PL science camera. Figure 15 shows a dark-subtracted averaged image after frame selection. We can already identify, in this data, three prominent absorption lines. We also observe that the flux distribution varies across traces, indicating the sensitivity of PL to the input scene or wavefront. We calculated the S/N for each minute of observation by dividing the average flux by the standard deviation of the flux, adjusted for the number of frames (i.e., the error on the flux measurement), per spectral channel over time. This approach yielded S/N values that varied around 200, primarily due to fluctuations in the injection efficiency.
@@ 4.2.1 Spectrum reconstruction
For each selected frame, we subtracted the dark+bias and extracted each PL output trace. The traces were then co-added for each frame, and then averaged over the entire observation sequence.
To calibrate the response of our instrument, we used a flat-field halogen lamp (approximately black body spectrum) positioned at the entrance of AO188 to uniformly illuminate the PL input FoV. By performing the same computation as was previously described, we derived a flat field spectrum, which we then calibrated using the theoretical black body slope corresponding to the Halogen lamp temperature (3020.3K). Subsequently, we utilized the resulting calibrated flat data to calibrate the Ikiiki reconstructed spectrum obtained with the PL. We further refined the computed spectrum by normalizing it to match a reference spectrum obtained from Alekseeva et al. (1996), represented by black dots in Fig. 16.
On the same graph, we show the resulting calibrated and adjusted spectrum of Ikiiki obtained with the PL. The latter closely matches the reference spectrum. Some differences can be noted, which may be attributed to the imperfect calibration of our data. Indeed, calibrating our reconstructed spectrum with the halogen lamp may potentially overlook chromatic perturbations caused by the atmosphere and/or various optical surfaces prior to AO188. Ideally, calibration should be conducted using a well-calibrated point source in the sky. The spectrum of Ikiiki exhibits three distinct absorption lines: the atmospheric oxygen A and B bands (around 761 nm and 687 nm, respectively), and H (at approximately 656 nm). The absorption in the two oxygen bands is due to oxygen molecules in Earth’s atmosphere and does not appear in the reference spectrum due to its terrestrial origin. At a spectral resolution of R3000, the absorption bands appear as a deep broad absorption line with a forest of narrower lines at longer wavelengths. The H absorption line originates from the cooler outer layers of the Ikiiki atmosphere, where neutral hydrogen atoms absorb photons emitted by the deeper, hotter layers at the energy level corresponding to the H transition.
@@ 4.2.2 Injection efficiency
In order to estimate the on-sky injection efficiency of the PL, we compared the acquired spectra (prior to flat-field calibration) to the reference spectro-photometric data of Ikiiki (Alekseeva et al., 1996). We summarize in Table 2 the throughput of elements from the entrance of the telescope down to the detector. We considered the detector quantum efficiency provided by the manufacturer (from to from 600 nm to 785 nm). We also modeled the throughput of the Earth’s atmosphere using the synthetic photometry Python package (synphot) maintained by the Space Telescope Science Institute (STScI). The airmass during the observation was 1.01, and induced an extinction ranging from 0.1 to 0.03 magnitude between 600 nm and 770 nm, leading to a throughput of on average about .
@@
Even under these difficult conditions, the collected data had a good S/N and enabled the faithful reconstruction of the Ikiiki spectrum.
@@ 4.3 Observation of ‘Aua
We observed ‘Aua for 53 minutes at a 200 Hz frame rate, providing a total of 636,000 frames. Similarly to the previous target, we performed a frame selection, since the conditions were not good. The best of frames were selected, dark+bias subtracted, and averaged to compute the image displayed in Fig. 19. Many spectral features can be identified on the averaged image. The computed S/N per minute was around 130, a sign that the observation conditions worsened compared to the previous target (where the S/N per minute was 200 in average).
After trace extraction and calibration with flat-field data, we co-added the resulting spectra to obtain the spectrum shown in Fig. 20. The spectrum shows multiple absorption spectral lines. The absorption lines detected on Ikiiki are present: oxygen bands A and B from the Earth’s atmosphere around 761 nm and 687 nm, respectively, and H absorption by the stellar atmosphere at about 656 nm. Many titanium oxide (TiO) lines are also present in the spectrum, with the strongest at about 705 nm. TiO lines in red supergiants are characteristic of convection cells within the star’s atmosphere (Young et al., 2003; Afram & Berdyugina, 2019), and are well known in the case of ’Aua as they are extensively exploited for image reconstruction using interferometry (Young et al., 2000; Levesque et al., 2005; Chiavassa et al., 2010). The ability to detect these (and the following) spectral lines, while conserving spatial information via the use of a PL opens the door to future use of differential phase interferometry (Akeson et al., 2000) to recover spatial features (such as stellar surface details) from PL observations. We can also spot the presence of vanadium oxide (VO) lines from about 730 nm to 750 nm. Similarly to TiO lines, VO lines are generated by convection cells in red supergiant stars, although they are less dominant than TiO. Finally, we identified the barium II (Ba II) line at about 649 nm (Koza, 2010). These lines can also be seen in data from the X-shooter Spectral Library (XSL) (Gonneau et al., 2020) taken at the VLT-UT2 using a medium-resolution spectrograph. This demonstration highlights the capabilities of the PL signal reconstruction to retrieve a target’s spectral features. The next step, not addressed in this demonstration paper, will be to exploit the spectral reconstruction to extract astronomical information about the observed target.
@@ 5 Conclusion
We have presented the integration, characterization, and on-sky demonstration of a PL at the 8.2 m Subaru Telescope. Photonic lanterns allow for high-efficiency SMF-based spectroscopy thanks to their unique design. Their MM input insures high coupling efficiency, compared to a single SMF, especially at visible wavelengths at which even ExAO systems struggle to reach high Strehl ratios. The SM outputs of PLs are perfectly suited for spectroscopy using compact and stable instruments.
@@
Depending on the optical injection setup focal ratio, the PL can provide various FoVs and injection efficiency performance depending on the desired science output. Our setup shows an optimal injection efficiency for a focal ratio of about 8, providing an injection efficiency of about over an FoV of 80 mas, with an efficiency peak of at 642 nm.
@@
We studied the relationship between the PL injection efficiency and Strehl ratio. Various turbulence profiles were simulated on the SCExAO DM, and the results were confronted to simulations in the case of an SMF. We have shown that PLs have a similar linear relationship between the injection efficiency and the Strehl to that of an SMF. However, PLs perform better in the presence of turbulence dominated by low-frequency aberrations.
@@
Finally, we tested the PL on-sky. Two stars were observed: a point source, Ikiiki ( Leo), and a red supergiant star, ‘Aua ( Ori). We were first able to reconstruct Ikiiki’s spectrum with high fidelity compared to its reference. Under median seeing conditions, the injection efficiency into the PL averaged to and reached up to at 642 nm, leading to an instrument throughput of up to . The best performance was obtained at 680 nm, where the injection efficiency into the PL averaged and reached up to , leading to an instrument throughput of up to . It is important to note that these figures likely fall short of demonstrating the full capabilities of our visible PL under more favorable conditions. For context, attempting a similar light injection directly into a single SMF under these conditions would likely yield inferior results. The observation of ‘Aua confirms our ability to reconstruct a complex spectrum from the PL data. We were able to retrieve multiple spectral features characteristic of this target.
@@
The on-sky performance should be increased under better atmospheric conditions, and even further with the soon-to-arrive upgrades of the Subaru AO facility DM (from 188 elements to 3000 elements) and visible wavefront sensor (Ahn et al., 2023). We also plan on upgrading the NPBS, picking up the light from SCExaO, and routing it to the PL with a R90/T10 BS. This should significantly increase the overall throughput of the instrument.
@@
This result shows that our setup can be used for high-throughput SMF-fed spectroscopy in the visible, especially for H spectro-imaging. The ongoing and near-future work aims to provide this capability to the community as a feature of the FIRST instrument. This demonstration constitutes the first on-sky acquisition of a spectrum using a PL fed by an ExAO system in the visible.
@@ Acknowledgements.
- Kim et al. (2024) Kim, Y. J., Fitzgerald, M. P., Lin, J., et al. 2024, The Astrophysical Journal, 964, 113
@@
- Tobin et al. (2024) Tobin, T. L., Currie, T., Li, Y., et al. 2024, The Astronomical Journal, 167, 205
@@
Appendix A Optical setup parameters for various numerical apertures of the beam injected into the PL
Table 3 provides optimized parameters of the PL injection bench for reference. Parameters are given for various focal ratios of the beam injected into the PL. Distances between the optical components are measured from the back surface of the first optical component to the front surface of the second optical component. For focal ratios of , , and , only two lenses are needed (L1 and L2). The back focal length provided in the table corresponds to the distance between the back surface of the last lens (L2 or L3, depending on the focal ratio considered) and the focal plane.